40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Mar 7 17:39:40 2012, keiko, Update, LSC, MICH noise budget on 5 Mar Mar5-MICHbudget.pngMar5-MICHbudget.png
    Reply  Wed Mar 7 20:53:13 2012, keiko, Update, LSC, MICH noise budget on 5 Mar Mar6-MICHbudget.png
    Reply  Thu Mar 8 00:57:48 2012, keiko, Update, LSC, MICH noise budget on Mar 5, Mar 6, and old Mar7MICHbudgettotal.png
Message ID: 6385     Entry time: Thu Mar 8 00:57:48 2012     In reply to: 6376
Author: keiko 
Type: Update 
Category: LSC 
Subject: MICH noise budget on Mar 5, Mar 6, and old 

Here is the recent two noise budgets of MICH, with the old measurement by Jenne. The most latest Mar 6 data is quite close to the old data, even better around 20-30 Hz. Probably some scattering source was improved?

Mar7MICHbudgettotal.png

ELOG V3.1.3-